Please help me solve this average questions:
1) Sabrina scored an average of 68 marks for 3 of her Mathematics tests. She wants to pull up her average score to 70 marks. How many marks does Sabrina need to score for her 4th test?

Answer by richwmiller(9143)

(3*68+x)/4=70
x=76
another way to look at it is
each of the first three missed 70 by 2
3*2=6
70 +6=76